# Exact Consistency Under Partial Views: Graph Colorability, Capacity, and Equality in Multi-Location Encodings

### 16 Mar 2026 v3

@@ VI Block Composition, Strong Powers, and Asymptotic Capacity
VI-A Motivation: Repeated Composition
What happens if we repeat the partial-view experiment $n$ times? If we compose $n$ independent copies of the same encoding system, does the ambiguity explode uncontrollably, or does it grow at a predictable rate?
The answer depends on the failure topology:
  • If the system were fully incoherent (confusability graph is a clique), every block would multiply the confusion. The ambiguity would grow without any exploitable structure.
  • Because our confusability graph has structure (as in the 4-cycle example of Section V-A), the ambiguity growth is controlled by that structure. Some state pairs remain distinguishable even across multiple blocks.
Under $n$-fold block composition, the confusability graph becomes its $n$-fold strong power:
$G_{n}\;=\;G^{\boxtimes n}.$
Independent sets in the strong power relate systematically to independent sets in the base graph, yielding a supermultiplicative growth law that converges to a well-defined asymptotic rate. The theorem below quantifies this rate.
VI-B Block Composition Law
@@
where $\alpha_{n}$ denotes the largest one-shot zero-error code size in the $n$-block composed system. This is the correct discrete precursor to Shannon-capacity analysis: the admissible code-size sequence is already supermultiplicative at the graph level.(L: MFT22)
More sharply, the block-level graph itself factors correctly. Confusability between concatenated $(m+n)$-block states is equivalent to adjacency in the strong product of the $m$-block and $n$-block confusability graphs, packaged formally as an explicit graph isomorphism. Thus the repeated-block system is not only bounded by strong-product arguments at the codebook level; it is literally a strong-product graph-power object for the same induced failure map.(L: MFT29-30, MFT38-39)
Conjunction, not disjunction.
Block confusability requires a single location tuple $\ell=(\ell_{1},\ell_{2})$ that works simultaneously for both blocks:
  • Block 1: observe location $\ell_{1}$, must match
  • Block 2: observe location $\ell_{2}$, must match
Two $(m+n)$-block states are confusable iff there exists one location tuple making observations match across all blocks. This is not “confusable in block 1 OR confusable in block 2” but rather $\exists$ location tuple, $\forall$ blocks: match.” That universal quantification over blocks is exactly the strong product condition: for each block, either equal or confusable (matching on some location), with at least one block strictly confusable.
Concrete example.
Consider two copies of the binary square. State $((0,0),(0,0))$ vs $((0,1),(1,1))$:
  • Block 1: $(0,0)\sim(0,1)$ in $C_{4}$ (agree on $x_{1}$)
  • Block 2: $(0,0)\not\sim(1,1)$ in $C_{4}$ (disagree on both coordinates)
An “OR-product” would mark these confusable because block 1 matches. But the actual model requires one location pair $(\ell_{1},\ell_{2})$ matching both blocks. Block 1 can match via $\ell_{1}=1$ (observing $x_{1}$). Block 2 cannot match: location 1 gives 0 vs $1$, location 2 gives 0 vs $1$. No location pair works, so these states are not confusable, exactly as the strong product predicts.
Iterating the same construction yields the derived $n$-block lower bound. If $\alpha_{1}$ denotes the one-shot maximum independent-set code size, then the mechanized block-power theorem gives
$\alpha_{1}^{n}\;\leq\;\alpha_{n},$
where $\alpha_{n}$ is the largest one-shot zero-error code size in the $n$-fold block-composed system. This is the first asymptotic-capacity scaffold in the extended model: repeated block composition already forces the expected exponential growth law before any Shannon-capacity or $\vartheta$-style upper theory is introduced. Once the architecture determines the one-shot failure graph, that same graph already controls the asymptotic lower law.(L: MFT21)
For notation, write $\Theta(G)$ for the Shannon capacity of a confusability graph $G$ in logarithmic units, and write $\vartheta_{\infty}(G)$ for the corresponding asymptotic Lovász-$\vartheta$ upper value. When the object is a view family $\mathcal{V}$ rather than an abstract graph, we write $\Theta(G_{\mathcal{V}})$ and $\vartheta_{\infty}(G_{\mathcal{V}})$ for the same quantities computed on the induced confusability graph $G_{\mathcal{V}}$.
These normalized block rates can be packaged into an explicit lower asymptotic capacity envelope
@@
implemented as an $\mathrm{ENNReal}$ supremum of the positive block-rate sequence. Formally, $\alpha_{n}$ is identified directly with the maximum finite independent-set size of the $n$-block confusability graph, the block graph is identified with the $n$-fold strong power of the one-shot confusability graph, and the same envelope is proved equal to the graph-side Shannon lower-capacity expression built from those strong powers.(L: MFT31-32, MFT35-36)
The asymptotic scaffold is also not limited to a lower envelope statement. On the graph side itself, strong-product lower bounds for independent-set cardinality induce monotonicity of normalized strong-power rates along repeated multiples. Specializing those generic graph theorems back to the model-generated confusability graph yields the repeated-block statement below.(L: MFT40-42)
Repeating a fixed $m$-block system $k$ times yields
$\alpha_{m}^{k}\;\leq\;\alpha_{km},$
and hence
$\frac{\log\alpha_{m}}{m}\;\leq\;\frac{\log\alpha_{km}}{km}.$
Thus rates along repeated block multiples are monotone from below toward the same capacity envelope. This is still a lower theory rather than a full Shannon-capacity theorem, but it is already a genuine asymptotic graph-rate statement for the induced failure topology.(L: MFT33)
Once the block confusability graph has been identified with strong powers of the one-shot graph, the asymptotic rate theory is exactly the classical Shannon-capacity framework specialized back to the model-generated graph family.
Theorem VI.1 (Asymptotic Shannon-capacity theorem).
@@
The argument rests on two observations: supermultiplicativity of independent-set sizes under strong product, and Fekete’s lemma.
Supermultiplicativity. Recall that the $n$-block confusability graph $G_{n}$ is the $n$-fold strong power $G^{\boxtimes n}$ of the one-shot graph $G$. For any two graphs $H$ and $K$, an independent set in $H\boxtimes K$ can be constructed from independent sets in $H$ and $K$ as follows. If $I_{H}\subseteq V(H)$ and $I_{K}\subseteq V(K)$ are independent, then the Cartesian product $I_{H}\times I_{K}\subseteq V(H)\times V(K)$ is independent in $H\boxtimes K$: if $(u,v)$ and $(u^{\prime},v^{\prime})$ are adjacent in the strong product, then either $u=u^{\prime}$ and $v\sim v^{\prime}$, or $u\sim u^{\prime}$ and $v=v^{\prime}$, or $u\sim u^{\prime}$ and $v\sim v^{\prime}$. In each case at least one of $u,u^{\prime}$ or $v,v^{\prime}$ is an edge within the respective independent set, which is impossible. Hence $\alpha(H\boxtimes K)\geq\alpha(H)\alpha(K)$. Iterating gives $\alpha_{m+n}=\alpha(G^{\boxtimes(m+n)})\geq\alpha(G^{\boxtimes m})\alpha(G^{\boxtimes n})=\alpha_{m}\alpha_{n}$, so the sequence $\{\alpha_{n}\}$ is supermultiplicative.
Fekete convergence. The inequality $\alpha_{m+n}\geq\alpha_{m}\alpha_{n}$ is equivalent to subadditivity of the sequence $\{-\log\alpha_{n}\}$. The trivial bound $\alpha_{n}\leq|V(G)|^{n}$ ensures that $\log\alpha_{n}\geq-n\log|V(G)|$, so the sequence is bounded below. By Fekete’s subadditivity lemma, the normalized limit
$\lim_{n\to\infty}\frac{\log\alpha_{n}}{n}=\sup_{n\geq 1}\frac{\log\alpha_{n}}{n}$
exists and equals the supremum of the finite block-rate sequence. This limit is the asymptotic Shannon capacity $\Theta(G)$ in logarithmic units.
The mechanized development packages the same argument with explicit sequence objects and convergence lemmas; the textual expansion above makes the independent-set product construction and Fekete application explicit for accessibility.  $\blacksquare$
@@
For the 4-cycle $C_{4}$ from Section V-A, the independence number is $\alpha(C_{4})=2$ (choose either diagonal). The strong power $C_{4}^{\boxtimes n}$ has $\alpha(C_{4}^{\boxtimes n})=2^{n}$, giving normalized rate $(\log 2^{n})/n=\log 2$. The asymptotic Shannon capacity is therefore $\Theta(C_{4})=\log 2$. This matches the exact tag-budget growth from the iterated binary-square family discussed earlier: the one-shot budget is $2$, and the $m$-block exact budget grows as $2^{m}$.
@@
This places the extended model directly in the classical zero-error graph-capacity setting. The difference from the classical channel picture is generative rather than formal: here the confusability graphs arise from deterministic view families and partial observations on latent tuples. We do not claim a new general Shannon-capacity theorem for arbitrary graphs; the new content is that this partial-view model reaches the classical graph machinery in a non-clique regime, so the architecture determines the graph once and that graph then determines both one-shot and asymptotic recoverability. The clique-fiber regime recovers the classical cluster-graph equality case, while the transitivity criterion identifies when that case is produced by the view-family model. The non-clique witness already shows that the model also generates genuinely nontrivial graph behavior.
VII Theta Upper Theory
This section supplies the asymptotic upper law for the same induced failure graph. The previous section showed that repeated composition turns view-induced ambiguity into strong powers and hence into an asymptotic recoverability rate. The question now is how large that rate can be. The first upper bound is complement-chromatic. The stronger package identifies the same one-shot upper invariant with standard orthonormal, primal-PSD, and dual-theta forms and then lifts those bounds to strong powers.
@@ VII Theta Upper Theory
and the asymptotic capacity is bounded above by $\log\chi(\overline{G})$.(L: MFT50-55)
At the one-shot level, the same upper invariant is matched with standard witness, orthonormal-representation, primal-PSD, and dual-theta formulations by explicit feasible-object equivalences. Applying these one-shot bounds to strong powers yields subadditive upper-rate sequences, so Fekete’s lemma gives asymptotic primal and dual upper values, each equal to the infimum of its finite power-rate bounds. The finite block rates are bounded above by either sequence, hence
@@
The same induced topology that governs exact recovery therefore also carries a classical asymptotic upper theory. The formal development proves these one-shot equivalences and the resulting asymptotic primal/dual upper laws.(L: MFT57-84)
Interpretationally, the complement-chromatic bound is the first combinatorial ceiling, while the Lovász-$\vartheta$ package supplies the sharper geometric and semidefinite upper theory attached to the same confusability graph. Once a deterministic partial-view architecture generates a failure graph, the standard zero-error upper hierarchy attaches to that graph without further model-specific modification.
@@ VIII Equality Characterization
Open question. A precise open problem is to characterize which upward‑closed agreement‑set families produce transitive confusability, and therefore trigger the cluster‑graph equality collapse. We have identified boundary facts: the binary square is a small witness where transitivity fails, while trivial single‑view families and the fiber‑coherent/meet‑witnessed families provably produce transitive confusability and hence the equality collapse. A complete classification of upward‑closed families that yield transitivity remains open and is an attractive direction for further work.
@@
At the witness level, this same collapse admits a local composition form: base confusability is the edge union of the single-view cluster graphs, and transitivity is equivalent to closure of those fixed-view witnesses under two-step composition through an intermediate state. This is the exact local criterion for the current equality mechanism: the collapse to a cluster graph is determined by how single-view witnesses compose, not by an opaque global coincidence of the full confusability graph. Thus the equality question is structural: it identifies when genuinely non-clique failure geometry collapses to exact cluster behavior.(L: MFT85-91, MFT100-102, MFT107-109, MFT136-137)
@@
The binary square of Section V-A does not have transitive confusability: $(0,0)\sim(0,1)$ and $(0,1)\sim(1,1)$, but $(0,0)\not\sim(1,1)$. The 4-cycle is not a cluster graph, so the transitivity-based equality collapse does not apply. In this special case the classical graph invariants still satisfy $\Theta(C_{4})=\vartheta_{\infty}(C_{4})=\log 2$, but that coincidence no longer comes from the cluster-collapse mechanism isolated in this section. The binary square therefore marks the boundary of the paper’s equality route: it is the first non-clique witness, yet not a transitive one.
@@
IX Affine Dual: Coordinate Matroid
The preceding sections developed the main graph-capacity and equality arc for partial views. This section turns to a second lens on the same model: which fact coordinates determine others across the realized state family? Under an affine restriction on the realized state family, that determination problem reduces to standard linear algebra and yields a representable matroid on fact indices. This matroid is the coordinate-side dual of the same structure that generates the confusability graph, and it provides computationally tractable upper bounds on confusability and capacity. This matroid perspective connects the problem to the rich literature on representable matroids in coding theory and combinatorial optimization, where matroid rank functions provide tractable bounds on code parameters.
Proposition IX.1 (Affine fact-matroid specialization).
Assume the realized state family is affine over a field, so that the valid latent tuples form an affine translate of a linear subspace of the ambient fact space. For any fact set $S$ and fact index $i$, semantic determination of $i$ by $S$ is equivalent to membership of the coordinate functional for $i$ in the linear span of the coordinate functionals indexed by $S$. Consequently the fact indices carry a representable matroid: a fact set is independent exactly when its coordinate functionals are linearly independent, and the minimal determining fact sets are exactly the bases, all of common cardinality equal to the rank.(L: AFM1-5)
@@ IX Affine Dual: Coordinate Matroid
The matroid and the confusability graph are complementary objects induced by the same realized-state model. Under the common specialization to finite affine families with coordinate-projection views, the matroid rank provides computationally tractable upper bounds on confusability and capacity. The tractability claim is representation-sensitive and should be read that way. We assume the affine family $A=a_{0}+V$ is given by an explicit linear presentation of $V$, for example a basis or generator matrix over $\mathbb{F}_{q}$, together with the explicit coordinate-subset view family. Under that input model, for each coordinate set $S$ the quantity $t(S)$ is exactly the finite-dimensional rank of the restricted coordinate map $\pi_{S}|_{V}$: the formalization identifies the span-based rank quantity with the finrank of the image of that explicit projection map.(L: AFM10-11) It also proves the basic size bounds $t(S)\leq|S|$ and $t(S)\leq\dim V$, equivalently $\dim(\operatorname{im}(\pi_{S}|_{V}))\leq\min\{|S|,\dim V\}$, and shows that $t(S)=|S|$ on linearly independent coordinate families while $t(S)$ reaches the full ambient determining rank exactly on determining sets.(L: AFM6-9, AFM12) Hence each $t(S)$ is a standard linear-algebraic rank quantity that can be computed by Gaussian elimination on an explicit presentation of $V$, and the bounds of Corollary IX.3 are then computable in polynomial time in the ambient dimension, the dimension of $V$, and the number of admissible views. Without an explicit linear presentation of $V$, no algorithmic tractability claim is intended. The utility of this matroid structure is immediate: $t(S)$ provides an explicit upper bound on the Shannon capacity of the induced confusability graph, bypassing the hardness of the general capacity computation. Computing Shannon capacity for general graphs is hard (the independence number is NP-hard, and even the Lovász-$\vartheta$ bound requires semidefinite programming). In contrast, once the affine family is explicitly presented, the relevant rank quantities reduce to standard linear algebra.
IX-A View-fiber dimensions and capacity

### 31 Mar 2026 v4

@@ V Block Composition, Strong Powers, and Asymptotic Capacity
V-A Motivation: Repeated Composition
Under $n$-fold block composition, the confusability graph becomes its $n$-fold strong power: $G_{n}=G^{\boxtimes n}$. Unlike the clique case where ambiguity explodes, structured graphs yield controlled growth via supermultiplicative independent-set behavior. The theorem below quantifies the resulting asymptotic rate.
V-B Block Composition Law
@@
where $\alpha_{n}$ denotes the largest one-shot zero-error code size in the $n$-block composed system. This is the correct discrete precursor to Shannon-capacity analysis: the admissible code-size sequence is already supermultiplicative at the graph level.464646: MFT22
More sharply, the block-level graph itself factors correctly. Confusability between concatenated $(m+n)$-block states is equivalent to adjacency in the strong product of the $m$-block and $n$-block confusability graphs, packaged formally as an explicit graph isomorphism. Thus the repeated-block system is not only bounded by strong-product arguments at the codebook level; it is literally a strong-product graph-power object for the same induced failure map.474747: MFT29-30, MFT38-39
Strong-product quantification.
Block confusability requires one location tuple working simultaneously for both blocks. Iterating yields $\alpha_{1}^{n}\leq\alpha_{n}$.484848: MFT21
For notation, $\Theta(G)$ denotes Shannon capacity of confusability graph $G$ and $\vartheta_{\infty}(G)$ the asymptotic Lovász-$\vartheta$ upper value. The lower asymptotic capacity envelope
@@
is identified with the maximum independent-set size of the $n$-block confusability graph.494949: MFT31-32, MFT35-36
Strong-product lower bounds give monotonicity: $\alpha_{m}^{k}\leq\alpha_{km}$ and $\frac{\log\alpha_{m}}{m}\leq\frac{\log\alpha_{km}}{km}$.505050: MFT33
Theorem V.1footnotemark: footnotetext: : MFT43-49, MFT56 (Asymptotic Shannon-capacity theorem).
@@
The $n$-block confusability graph is $G_{n}=G^{\boxtimes n}$. For any graphs $H,K$, the Cartesian product of independent sets $I_{H}\times I_{K}$ is independent in $H\boxtimes K$, giving $\alpha(H\boxtimes K)\geq\alpha(H)\alpha(K)$. Iterating yields supermultiplicativity $\alpha_{m+n}\geq\alpha_{m}\alpha_{n}$, equivalent to subadditivity of $\{-\log\alpha_{n}\}$. Boundedness follows from $\alpha_{n}\leq|V(G)|^{n}$. Fekete’s lemma [17] gives
$\lim_{n\to\infty}\frac{\log\alpha_{n}}{n}=\sup_{n\geq 1}\frac{\log\alpha_{n}}{n},$
the asymptotic Shannon capacity $\Theta(G)$.  $\blacksquare$
@@
For the 4-cycle $C_{4}$ from Section IV-A, $\alpha(C_{4})=2$ and $\alpha(C_{4}^{\boxtimes n})=2^{n}$, giving normalized rate $\log 2$ and asymptotic Shannon capacity $\Theta(C_{4})=\log 2$. This matches the iterated binary-square family: one-shot budget is $2$, $m$-block budget grows as $2^{m}$.
@@
This places the extended model in the classical zero-error graph-capacity setting. Confusability graphs arise from deterministic view families on latent tuples; the architecture determines the graph, which then determines recoverability. The clique-fiber regime recovers the cluster-graph equality case via transitivity of confusability.
VI Theta Upper Theory
This section supplies the asymptotic upper law for the same induced failure graph. The previous section showed that repeated composition turns view-induced ambiguity into strong powers and hence into an asymptotic recoverability rate. The question now is how large that rate can be. The first upper bound is complement-chromatic. The stronger package identifies the same one-shot upper invariant with standard orthonormal, primal-PSD, and dual-theta forms and then lifts those bounds to strong powers [3, 18, 19, 20].
@@ VI Theta Upper Theory
and the asymptotic capacity is bounded above by $\log\chi(\overline{G})$.525252: MFT50-55
At the one-shot level, the same upper invariant is matched with standard witness, orthonormal-representation, primal-PSD, and dual-theta formulations by explicit feasible-object equivalences. Applying these one-shot bounds to strong powers yields subadditive upper-rate sequences, so Fekete’s lemma [17] gives asymptotic primal and dual upper values, each equal to the infimum of its finite power-rate bounds. The finite block rates are bounded above by either sequence, hence
@@
The same induced topology that governs exact recovery therefore also carries a classical asymptotic upper theory. The formal development proves these one-shot equivalences and the resulting asymptotic primal/dual upper laws.535353: MFT57-84
Interpretationally, the complement-chromatic bound is the first combinatorial ceiling, while the Lovász-$\vartheta$ package supplies the sharper geometric and semidefinite upper theory attached to the same confusability graph. Once a deterministic partial-view architecture generates a failure graph, the standard zero-error upper hierarchy attaches to that graph without further model-specific modification [3, 21, 22, 23].
@@ VII Equality Characterization
Open problem. A precise open problem is to characterize which upward‑closed agreement‑set families produce transitive confusability, and therefore trigger the cluster‑graph equality collapse. We have identified boundary facts: the binary square is a small witness where transitivity fails, while trivial single‑view families and the fiber‑coherent/meet‑witnessed families provably produce transitive confusability and hence the equality collapse. A complete classification of upward‑closed families that yield transitivity remains open and is an attractive direction for further work.
@@
At the witness level, this same collapse admits a local composition form: base confusability is the edge union of the single-view cluster graphs, and transitivity is equivalent to closure of those fixed-view witnesses under two-step composition through an intermediate state. This is the exact local criterion for the current equality mechanism: the collapse to a cluster graph is determined by how single-view witnesses compose, not by an opaque global coincidence of the full confusability graph. Thus the equality question is structural: it identifies when genuinely non-clique failure geometry collapses to exact cluster behavior.545454: MFT85-91, MFT100-102, MFT107-109, MFT136-137
@@
The binary square of Section IV-A does not have transitive confusability: $(0,0)\sim(0,1)$ and $(0,1)\sim(1,1)$, but $(0,0)\not\sim(1,1)$. The 4-cycle is not a cluster graph, so the transitivity-based equality collapse does not apply. In this special case the classical graph invariants still satisfy $\Theta(C_{4})=\vartheta_{\infty}(C_{4})=\log 2$, but that coincidence no longer comes from the cluster-collapse mechanism isolated in this section. The binary square therefore marks the boundary of the paper’s equality route: it is the first non-clique witness, yet not a transitive one.
@@
VIII Affine Dual: Coordinate Matroid
The preceding sections developed the main graph-capacity and equality arc for partial views. Computing the resulting graph quantities is hard in general. This section isolates a regime in which the coordinate side of the model becomes tractable: when the realized state family is affine, semantic determination reduces to standard linear algebra, and the resulting upper bounds can be computed by matrix rank rather than by solving a general graph-capacity problem.
The resulting object is a representable matroid on fact indices. It is the coordinate-side dual of the same realized-state structure that generates the confusability graph, and it provides computationally tractable upper bounds on confusability and capacity. The tractability claim is representation-sensitive: throughout this section, the affine family $A=a_{0}+V$ is assumed to be given by an explicit linear presentation of $V$, for example a basis or generator matrix over $\mathbb{F}_{q}$, together with the explicit coordinate-subset view family. Under that input model, the relevant quantity is just the rank of a restricted coordinate map and is computable by Gaussian elimination.
Warm-up example.
Before stating the general proposition, consider the binary square from Section IV-A. Here the realized state family is the full affine space $A=\mathbb{F}_{2}^{2}$, so we may take $a_{0}=(0,0)$ and direction space $V=\mathbb{F}_{2}^{2}$ with generator matrix
$G=\begin{bmatrix}1&0\\ \ 0&1\end{bmatrix}$.
The coordinate functionals are $e_{1}=[1\ \ 0]$ and $e_{2}=[0\ \ 1]$. A single-coordinate view, say $S=\{1\}$, has rank $t(S)=1$: it captures one of the two available dimensions, so it reveals exactly half of the informational dimension of the state space. This is the finite-affine meaning of the later matroid rank bound: the coordinate-side quantity $t(S)$ measures how much of the state survives the view, and in the affine regime it is obtained by ordinary matrix rank.
Proposition VIII.1 (Affine fact-matroid specialization).
Assume the realized state family is affine over a field, so that the valid latent tuples form an affine translate of a linear subspace of the ambient fact space. For any fact set $S$ and fact index $i$, semantic determination of $i$ by $S$ is equivalent to membership of the coordinate functional for $i$ in the linear span of the coordinate functionals indexed by $S$. Consequently the fact indices carry a representable matroid: a fact set is independent exactly when its coordinate functionals are linearly independent, and the minimal determining fact sets are exactly the bases, all of common cardinality equal to the rank.555555: AFM1-5
@@ VIII Affine Dual: Coordinate Matroid
The matroid and the confusability graph are complementary objects induced by the same realized-state model. Under the common specialization to finite affine families with coordinate-projection views, the matroid rank provides computationally tractable upper bounds on confusability and capacity. For each coordinate set $S$ the quantity $t(S)$ is the rank of the restricted coordinate map $\pi_{S}|_{V}$.565656: AFM10-11 The formalization proves the basic size bounds $t(S)\leq|S|$ and $t(S)\leq\dim V$, and shows that $t(S)=|S|$ on linearly independent coordinate families while $t(S)$ reaches the full ambient determining rank exactly on determining sets.575757: AFM6-9, AFM12 Hence each $t(S)$ is computable by Gaussian elimination on an explicit presentation of $V$, and the bounds of Corollary VIII.3 are computable in polynomial time. The utility is immediate: $t(S)$ provides an explicit upper bound on the Shannon capacity of the induced confusability graph, bypassing the hardness of the general capacity computation. Computing Shannon capacity for general graphs is hard (the independence number is NP-hard, and even the Lovász-$\vartheta$ bound requires semidefinite programming [24, 18]). In contrast, once the affine family is explicitly presented, the relevant rank quantities reduce to standard linear algebra.
VIII-A View-fiber dimensions and capacity